On summer day, the solar irradiation in Athens, Georgia provides 735 W/m2 at noon time with ambient temperature 26C. Find the temperature derate factor for the Hanwha 300W single crystal silicon solar module at noon. The NOCT temperature is 45C and the temperature coefficient of Pmax is -0.39% per C
Answer #1

Solar irradiance (SI) = power per unit area received from the Sun = 735 W/m2 = 0.735 kW/m2

Ambient temperature Ta= 26oC

Nominal Operating Cell Temperature TNOCT = 45oC

Temperature coefficient of Pmax = -0.39% per oC

Temperature of solar cell can be determined from following expression:

Tcell = Ta + 1.25 *SI* ( TNOCT  - 20) = (26 + 1.25*735*(45-20))oC = 48.97oC [SI has to be expressed in kW/m2 to use in this expression]

Assuming the rated power of the module is measured at 25 degrees Celsius (per standard practice),

power derate factor at temperature Tcell is given by

Pmax x Temperature coefficient x (Tcell - 25) = 300 W x (-0.39%) /oC x (48.97 - 25)oC = -28.045 W

i.e. The module will produce 28.045 W less power than the rated power for the given conditions.